OpenStudy (he66666):

linear algebra (cross product in r^3) find parametric equations of the line passing through (3, -1, -3) and perpendicular to the line passing through (3, -2, 4) and (0, 3, 5). answer: x=3, y=-1+t, z=-3-5t How do you solve this? I don't know how to get the direction vectors(?). All I know is that the direction vector of the given line is (-3, 5, 1) and you're supposed to find the perpendicular vector for that...

OpenStudy (kinggeorge):

Assuming that the line you're supposed to find is also supposed to intersect with the line you were given, I have a method, but it isn't too pretty.

OpenStudy (kinggeorge):

First, the direction of the first line you were given is <-3, 5, 1>. Note that any vector perpendicular to this line will lie in some plane that is perpendicular to that line. Hence, we are looking for a plane given by the equation \(-3x+5y+z+d=0\) such that this plane contains the point (3, -1, -3). Do you follow so far?

OpenStudy (he66666):

Yes. But we haven't dealt with planes. We did in high school, but I'm not sure if we can solve using planes now :/

OpenStudy (kinggeorge):

I'm not sure how to do this without planes. I'll finish what I have so far and see where we get though.

OpenStudy (kinggeorge):

If we use the plane, we get that our equation for the plane must be \[-3(3)+5(-1)+3(-3)+d=0\]Solving for \(d\), we get that \(d=23\). So our plane equation is \[-3x+5y+z+23=0\]Now we need to find where this plane intersects the first line given to us.

OpenStudy (kinggeorge):

We have that \[x=3-3t\]\[y=-2+5t\]\[z=4+t\]Substituting these into our plane equation, we get that \(t=-8/35\)

OpenStudy (kinggeorge):

Now we plug this back into the equations for \(x, y, z\) to get that it intersects at the point \((129/35,\;-22/7,\; 132/35)\)

OpenStudy (kinggeorge):

Using this as a second point, and \((3, -1, -3)\) as the first, we get a vector in the direction of \(<24/25, -29/7, 237/35>\) I seem to have done something wrong :(

OpenStudy (he66666):

Wow those are big numbers.. I think I may have got it, but I'm not sure. Since the direction vector is (-3, 5, 1), we can use the dot product (Assuming the perpendicular vector is (a, b, c): -3a+5b+c=0. And then can't we just plug in random numbers? So if we plug in 0 and 1 for a and b respectively, we automatically get -5 as c.

OpenStudy (kinggeorge):

That would probably work better than what I tried to do. Although you probably won't necessarily get an intersection point. I guess you don't need that.

OpenStudy (kinggeorge):

However, their answer for the line doesn't intersect the original line, so whatever.

OpenStudy (he66666):

Oh I see. Thanks for your help KingGeorge! :)

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (kinggeorge):

I think I see where I went wrong though. When solving for \(d\), I used the equation\[-3x+5y-3z+d=0\]When I should have used\[-3x+5y+z+d=0\]Derp :(

OpenStudy (he66666):

Ah I see haha. So I guess it should work using the method!

OpenStudy (kinggeorge):

Let's see where I get with the new numbers. Give me a minute.

OpenStudy (kinggeorge):

I get a vector in the direction of \(<6/35,\;-9/7,\;243/35>\) which still isn't pretty. However, when tested with the dot product, this is perpendicular to the original line.

OpenStudy (kinggeorge):

So an alternative answer (if you multiply through by 35) would be \[\vec r=\left(\begin{matrix}3\\-1\\-3\end{matrix}\right)+t\left(\begin{matrix}6\\-45\\243\end{matrix}\right)\]Which is a vector perpendicular to the given line, passes through the specified point, and also passes through the given line.

OpenStudy (he66666):

I guess there are a few possible answers, not just one? Thanks again :)

OpenStudy (kinggeorge):

You're welcome =D